# The cognitive triple-slit experiment

### v1

@@ 4 The cognitive triple-slit experiment
In this section, we explain how our cognitive experiment was conducted in practice. For simplicity, we only considered a one-dimensional detection screen. Note that the typical interference fringes are a consequence of the slits’ geometry, in the sense that it is their height, and the fact that they are parallel to each other, that are responsible for the fringe appearance, while it is their distance from each other, and width, that are responsible for the alternation between high and low impact zones and diffraction phenomena. For the purposes of our experiment, however, it was not necessary to investigate the vertical dimension of the slits. More precisely, our detection screen consisted of $n=8d+3$ elementary cells of unit length, where $d=5$ corresponds to the distance between the slits (the number of cells separating them) in the slit plate. The latter was placed at a distance of $L=12$ cells from the detection screen, which was formed of $n=43$ cells, with discrete positions $x=1,\dots,n$. The slits were chosen to be of unit width and placed in correspondence with cells $n=16,22,28$ (see Figures 1 and 2).
@@
A number $N=213$ of subjects participated in our interactive online experience in which they were asked to complete seven tasks in response to specific questions (corresponding to three one-slit, three double-slit and one triple-slit experiments). They were chosen at random from colleagues and friends and also recruited via a YouTube video on one of the authors’ channel. They were all exposed to the same questions and experimental conditions (repeated measures design). Also, the order of the one-slit and two-slit experiments were randomized for each subject, to minimize order effects in data collection. The number of participants recruited from March 6 to April 6, 2025, was actually $221$, but we excluded those who reported that the instructions were too difficult to understand or who took too much time (more than $30$ minutes) or too little time (less than $3$ minutes) to complete the experiment (the average time was $9.27$ minutes).
@@
More precisely, after agreeing to participate, each participant had to read an introductory text explaining the context of the experiment. Note that the introductory text, and subsequent explanatory texts, did not mention that it was a cognitive simulation of a quantum physics experiment, so as not to condition their choices on the basis of possible prior knowledge of the double and possibly triple-slit experiments. For this reason, the different situations were presented to them in purely imaginative terms, as a description of an animal looking for food to eat (see below). In addition to the introductory text, each of the tasks was preceded by an explanation of the criterion by which they had to select a certain number of points (cells) on the detection screen. Basically, in each task participants had to select $7$ cells in sequence. The first selection corresponded to what was considered the best cell to answer the question asked, receiving (during analysis) a score of $7$ points. The second selected cell, which could not be the same because once a cell was selected the system disabled it and could not be selected a second time, received a score of $6$ points, and so on, until the seventh selected cell, which received only $1$ point. In other words, respondents answered the questions using a rank-ordering procedure with a $7$-point score assignment applied during analysis. This is of course not exactly what happens in a physics laboratory, where a single impact at a time is actualized by the measuring instrument. However, if each subject were asked to select only one cell per task, the number of data collected would have been insufficient. Note also that if we would have assigned the same score to all their choices, the patterns we obtained would have remained essentially the same, only slightly less pronounced.
@@
Figure 1: The large rectangular area that was presented to the participants, here in the situation with three openings.
@@
Upon completion of the cognitive test, each of the $N=213$ admitted participants distributed $\sum_{i=1}^{7}i=28$ points to $7$ different cells for each of the seven tasks. Thus, a total of $N_{\rm tot}=28N=5964$ datapoints were assigned to the $n=43$ cells of the detection screen. If $N(x|1)$, $N(x|2)$ and $N(x|3)$ denote the number of points assigned to cell $x$, when only slit-$1$, slit-$2$ and slit-$3$ are open, respectively, dividing them by $N_{\rm tot}$ one obtains the corresponding one-slit experimental probabilities $P(x|1)$, $P(x|2)$ and $P(x|3)$. One does the same with the numbers $N(x|12)$, $N(x|23)$ and $N(x|13)$, when only ‘slit-$1$ and $2$’, ‘slit-$2$ and $3$’, and ‘slit-$1$ and $3$’ are open, respectively, and with the number $N(x|123)$, when all three slits are open, obtaining in this way the two-slit experimental probabilities $P(x|12)$, $P(x|23)$ and $P(x|13)$, and the three-slit probability $P(x|123)$, $x=1,\dots,n$. All these probabilities are reported in Table 1, with the Sorkin parameter $\epsilon(x|123)$ and its normalized version $\kappa(x|123)=P_{\rm max}^{-1}\,\epsilon(x|123)$, $P_{\rm max}=\sup_{x}P(x|123)$. Note that assuming a binomial distribution, the maximum value for the standard error is ${\rm SE}(19)\approx 0.00382$, where $x=19$ corresponds to the maximum probability $P_{\rm max}=P(19|123)\approx 0.0961$. If we consider a confidence level of $95\%$, the $z$-value is $1.96$ and the confidence interval becomes: $P(19|123)\approx 0.0961\pm 1.96\cdot 0.00382=0.0961\pm 0.0075$.
@@ 5 Analysis of the experimental data
Having summarized the experimental values reported in the physics literature, we can proceed to analyze our cognitive data, which we have represented graphically in Figures 2 and 3. Those for the three one-slit experiments, described in the left column of Figure 2, do not present any problems of interpretation. The question posed to the participants did not represent a situation of doubt, as there was certainty about the slit (the opening) from which the quantum entity (the animal) would emerge, and this explains the single peak forming in front of each of the slits. To estimate a possible effective wavelength $\lambda$ involved, we note that the peaks have a width of approximately $\Delta x\approx 10$. If we use the Fraunhofer diffraction formula $\Delta x\approx 2{\lambda L\over a}$, where $a$ is the width of the slit, since we have $a=1$ and $L=12$, the effective wavelength is $\lambda\approx a{\Delta x\over 2L}\approx{10\over 24}\approx 0.4$, i.e., approximately half a cell.
@@
Let us now consider the data obtained for the experiments with two slits, described in the right column of Figure 2. All three experiments show a clear central peak. This was of course to be expected, as the question posed attempted to maximize doubt about the slit (the opening) from which the quantum entity (the animal) emerged, and a central impact does indeed reflect maximum uncertainty. Another element common to the three two-slit probabilities in Figure 2 is the fact that the participants also placed many of their points at the extremes of the detection screen. The idea here is that by moving as far away as possible from the two slits, it becomes more difficult to discriminate between them, from a spatial perspective, so in some way a new situation of (local) maximum doubt is obtained about the slit from which the entity emerged. However, we also observe that the maximum is not always found at the outermost cells; this is the case, for example, in the graph for $P(x|12)$. One participant commented that choosing the most distant cells would have favored the closest slit. The logic behind this thought could be the following. If propagation from the slit to the point of impact has a cost proportional to the distance traveled, and resources are limited, then an entity emerging from slit-$1$ will be able to go further to the left than an identical entity (with same resources) emerging from slit-$2$. This suggests that if we had used a larger screen, a wide secondary fringe could have possible formed, at a distance not necessarily corresponding to its extreme sides, and this could be one of the cognitive mechanism at the origin of secondary regions of probability overextension, in addition to the central one.
@@
There is a further aspect that is particularly evident in the graph of $P(x|13)$, where the two slits are at a distance twice as far apart compared to the other two experiments. There are two additional secondary peaks in correspondence with the positions of the slits. These can also be glimpsed in the graph of $P(x|23)$. This choice doesn’t seem logical, but it was favored by a sufficient number of participants to make it emerge in the structure. The positions corresponding to the slits seem to be attractive in some way, although they tend to minimize rather than maximize doubt.
If we apply the well-known formula $\lambda\approx{d^{2}\over L}$, giving the approximate distance at which a second maximum for the probability forms in the far field approximation, [^2] However, as we will discuss later in this article, this approximation does not adequately describe our experiment, given that the nullity of Sorkin’s parameter is strongly violated. considering $d=6$ and $L=12$, we find an effective wavelength $\lambda\approx 3$, which is clearly not compatible with the previously determined $\lambda\approx 0.4$ describing an effective diffraction mechanism. This means that there is no way to associate a single effective wavelength to our cognitive experiments, and this would be an important difference with the situation in physics laboratories (see also the discussion in Section 6).
@@
Turning now to the situation with three slits (the central graph at the bottom of Figure 2), we always have accumulation points at the ends of the screen, and secondary peaks at each of the three slits, plus two major peaks at the midpoint between the slits. We also observe a strong asymmetry between left and right as far as the two midpoints are concerned, the midpoint between slits $1$ and $2$ being chosen much more frequently than the midpoint between slits $2$ and $3$. The left side of the screen was also favored for peaks in correspondence with slits $1$ and $2$ and for the accumulation points at the ends of the screen, confirming a general bias favoring the left side of the screen. The origin of this asymmetry remains unclear. It may be attributed to the fact that the participants are from a left-to-right reading culture, or it could be due to the population’s asymmetry between left-handed and right-handed individuals, but these are hypotheses that would require validation.
@@
When comparing our graph with those typically obtained in physics laboratories, the most striking difference is perhaps the fact that the central fringe is not the one of maximum intensity. A possible overall explanation for the existence of the two main peaks between slits $1$ and $2$, and between slits $2$ and $3$, for the secondary peak in correspondence with slit $3$, and for the two regions of accumulation of lesser intensity at the extremes of the screen, may be the following. When choosing a point $x$ on the screen, which should maximize the doubt about the slit from which the entity emerged, a participant will tend to associate with each slit a probability $\rho_{i}(x)$, $i=1,2,3$, inversely proportional to the distance $L_{i}$ between the slit and the point $x$ in question. In other words, the shorter the distance between the slit and point $x$, the more likely it is that the entity that produced that point emerged from that slit. Considering then the maximum $\rho(x)={\rm max}\{\rho_{1}(x),\rho_{2}(x),\rho_{3}(x)\}$, a participant trying to identify a point $x$ on the screen that maximizes the doubt about the slit from which the entity emerged will intuitively identify a position $x$ that minimizes $\rho(x)$. So, if $\overline{\rho}(x)=1-\rho(x)$ and $B=\sum_{x}\overline{\rho}(x)$, we can define $P_{\rm th}(x)=B^{-1}\overline{\rho}(x)$ as the “theoretical” probability for selecting a point $x$ when using a reasoning based on competing distances. In the left graph of Figure 3, we have plotted $P_{\rm th}(x)$ together with $P(x|123)$. We see that $P_{\rm th}(x)$ has its two main peaks precisely between slits $1$ and $2$, and between slits $2$ and $3$. This would explain the participants’ choice to favor those positions. Furthermore, the value of $P_{\rm th}(x)$ is higher between the two slits than outside them, and this would explain the secondary choice of the central position. Finally, $P_{\rm th}(x)$ shows accumulation points at the ends of the screen, which is also in line with the participants’ choices.
@@
Figure 3: (color online) Left: Plot of the triple-slit probability $P(x|123)$ (red color) together with the “theoretical” probability $P_{\rm th}(x)$ (black color) describing a reasoning based on the evaluation of distances between slits and impact points. To make the variations of $P_{\rm th}(x)$ easier to read, it has been drawn using a distance $L=3$ (instead of $L=12$). As $L$ decreases (near-field limit), $P_{\rm th}(x)$ increases at the edges of the screen, exceeding its value at the center. Right: Plot of the normalized Sorkin parameter $\kappa(x|123)$ (blue color). The probability $P(x|123)$, multiplied by a factor of $10$, is also plotted (red color), to give a clearer understanding of the modulation in the plot. A continuous approximation of the discrete functions was used to enhance interpretability in the visualization.
@@
Let us now consider the normalized Sorkin parameter $\kappa(x|123)$. Its maximum value is for $x=16$, i.e., in correspondence with slit-$1$: $\kappa(16|123)=0.549$; see Table 1 and the right graph of Figure 3. The values we obtained are much higher than those typically obtained in physics laboratories. This reveals the presence of significant deviations of the probability $P(x|123)$ both from the classical situation and from the idealized quantum situation, when the principle of superposition can be applied naively. In other words, in their decision-making, participants were also guided by considerations that led to an overall assessment of the situation with three slits, not reducible to a combination of reasonings on two slits at a time. In other words, they generated significant third-order interference contributions, which in physics can only be obtained in near-field situations, i.e., when the detection screen is not located far from the slits and/or slits are very close together (distances on the order of the wavelength or less), so that non-radiative modes, such as evanescent waves, can easily mediate coupling between the slits and the detector, and/or between the slits. It could be said that since the human cognitive instrument apparently uses different effective wavelengths at the same time, and wavelengths determine the scale of near-field dominance, in our double and triple-slit cognitive experiments we are always in regimes where near-field effects are important, hence the strong violation of the nullity of Sorkin parameter.

### v2

@@ 4 The cognitive triple-slit experiment
In this section, we explain how our cognitive experiment was conducted in practice. For simplicity, we only considered a one-dimensional detection screen. Note that the typical interference fringes are a consequence of the slits’ geometry, in the sense that it is their height, and the fact that they are parallel to each other, that are responsible for the fringe appearance, while it is their distance from each other, and width, that are responsible for the alternation between high and low impact zones and diffraction phenomena. For the purposes of our experiment, however, it was not necessary to investigate the vertical dimension of the slits. More precisely, our detection screen consisted of $n=8d+3$ elementary cells of unit length, where $d=5$ corresponds to the distance between the slits (the number of cells separating them) in the slit plate. The latter was placed at a distance of $L=12$ cells from the detection screen, which was formed of $n=43$ cells, with discrete positions $x=1,\dots,n$. The slits were chosen to be of unit width and placed in correspondence with cells with numbers $n=16,22,28$ (see Figures 4 and 6).
@@
A number $N=213$ of subjects participated in our interactive online experience in which they were asked to complete seven tasks in response to specific questions (corresponding to three one-slit, three double-slit and one triple-slit experiments). They were chosen at random from colleagues and friends and also recruited via a YouTube video on one of the authors’ channel. They were all exposed to the same questions and experimental conditions (repeated measures design). Also, the order of the one-slit and two-slit experiments were randomized for each subject, to minimize order effects in data collection. The number of participants, recruited from March 6 to April 6, 2025, was actually $221$, but we excluded those who reported that the instructions were too difficult to understand or who took too much time (more than $30$ minutes) or too little time (less than $3$ minutes) to complete the experiment (the average time was $9.27$ minutes).
@@
More precisely, after agreeing to participate, each participant had to read an introductory text explaining the context of the experiment. Note that the introductory text, and subsequent explanatory texts, did not mention that it was a cognitive simulation of a quantum physics experiment, so as not to condition their choices on the basis of possible prior knowledge of the double-slit and possibly triple-slit experiments. For this reason, the different situations were presented to them in purely imaginative terms, as a description of an animal looking for food to eat (see below). In addition to the introductory text, each of the tasks was preceded by an explanation of the criterion by which they had to select a certain number of points (cells) on the detection screen. Basically, in each task participants had to select $7$ cells in sequence. The first selection corresponded to what was considered the best cell to answer the question asked, receiving (during analysis) a score of $7$ points. The second selected cell, which could not be the same because once a cell was selected the system disabled it and could not be selected a second time, received a score of $6$ points, and so on, until the seventh selected cell, which received only $1$ point. In other words, respondents answered the questions using a rank-ordering procedure, with scores from $7$ to $1$ assigned during analysis. This is of course not exactly what happens in a physics laboratory, where the measuring instrument records only one impact at a time. However, if each subject had been asked to select only one cell per task, the number of data collected might have been insufficient. However, the patterns highlighted in this work remain identifiable regardless of the score that is attributed to the sequential choices of the participants, as we will explain in Section 5.
@@
Figure 4: (color online) The large rectangular area that was presented to the participants, with the row of $43$ identical portions of food, here in the situation with three openings.
@@
Upon completion of the cognitive test, each of the $N=213$ admitted participants distributed $\sum_{i=1}^{7}i=28$ points to $7$ different cells for each of the seven tasks. Thus, a total of $N_{\rm tot}=28N=5964$ datapoints were assigned to the $n=43$ cells of the detection screen. If $N(x|1)$, $N(x|2)$ and $N(x|3)$ denote the number of points assigned to cell $x$, when only slit-$1$, slit-$2$ and slit-$3$ are open, respectively, dividing them by $N_{\rm tot}$ one obtains the corresponding one-slit experimental probabilities $P(x|1)$, $P(x|2)$ and $P(x|3)$. One does the same with the numbers $N(x|12)$, $N(x|23)$ and $N(x|13)$, when only ‘slit-$1$ and $2$’, ‘slit-$2$ and $3$’, and ‘slit-$1$ and $3$’ are open, respectively, and with the number $N(x|123)$, when all three slits are open, obtaining in this way the two-slit experimental probabilities $P(x|12)$, $P(x|23)$ and $P(x|13)$, and the three-slit probability $P(x|123)$, $x=1,\dots,n$. All these probabilities are reported in Table 1, with the Sorkin parameter $\epsilon(x|123)$ and its normalized version $\kappa(x|123)=P_{\rm max}^{-1}\,\epsilon(x|123)$, $P_{\rm max}=\sup_{x}P(x|123)$.
Note that assuming a binomial distribution, the maximum value for the standard error is ${\rm SE}(19)\approx 0.00382$, where $x=19$ corresponds to the maximum probability $P_{\rm max}=P(19|123)\approx 0.0961$. If we consider a confidence level of $95\%$, the $z$-value is $1.96$ and the confidence interval becomes: $P(19|123)\approx 0.0961\pm 1.96\cdot 0.00382=0.0961\pm 0.0075$; see also Figure 5.
Figure 5: (color online) The evolution of the three-slit probabilities $P(x|123)$, $x=1,\dots,43$, as a function of the number of participants who progressively took part in the cognitive experiment, from $N=1$ to $N=213$. We observe that the curves stabilize as the number of participants increases, consistent with the law of large numbers.
@@ 5 Analysis of the experimental data
Having summarized the experimental values reported in the physics literature, we can proceed to analyze our cognitive data, which we have represented graphically in Figures 6 and 9. Those for the three one-slit experiments, described in the left column of Figure 6, do not present any problems of interpretation. The question posed to the participants did not represent a situation of doubt, as there was certainty about the slit (the opening) from which the quantum entity (the animal) would emerge, and this explains the single peak forming in front of each of the slits.
@@
Considering the data for the experiments with two slits, described in the right column of Figure 6, we observe that all three experiments show a clear central peak. This was of course to be expected, as the question posed attempted to maximize doubt about the slit (the opening) from which the quantum entity (the animal) emerged, and a central impact does indeed reflect maximum uncertainty. Another element common to the three two-slit probabilities in Figure 6 is the fact that participants also placed many of their points at the extremes of the detection screen. The idea here is that by moving as far away as possible from the two slits, it becomes more difficult to discriminate between them, from a spatial perspective, so in some way a new situation of (local) maximum doubt is obtained about the slit from which the entity emerged. However, we also observe that the maximum is not always found at the outermost cells; this is the case, for example, in the graph for $P(x|12)$. One participant commented that choosing the most distant cells would have favored the closest slit. The logic behind this thought could be the following. If propagation from the slit to the point of impact has a cost proportional to the distance traveled, and resources are limited, then an entity emerging from slit-$1$ will be able to go further to the left than an identical entity (with same resources) emerging from slit-$2$. This suggests that if we had used a larger screen, a wide secondary fringe could have possible formed, at a distance not necessarily corresponding to its extreme sides, and this could be one of the cognitive mechanism at the origin of secondary regions of probability overextension, in addition to the central one.
@@
There is a further aspect that is particularly evident in the graph of $P(x|13)$, where the two slits are at a distance twice as far apart compared to the other two experiments. There are two additional secondary peaks in correspondence with the positions of the slits. These can also be glimpsed in the graph of $P(x|23)$. This choice doesn’t seem rational, but it was favored by a sufficient number of participants to make it emerge in the structure. The positions corresponding to the slits seem to be attractive in some way, although they tend to minimize rather than maximize doubt. Therefore, pure logical-rational thought processes were not the only ones involved in the decision-making process..
@@
Turning now to the situation with three slits (the central graph at the bottom of Figure 6), we always have accumulation points at the ends of the screen, and secondary peaks at each of the three slits, plus two major peaks at the midpoint between the slits. We also observe a strong asymmetry between left and right as far as the two midpoints are concerned, the midpoint between slits $1$ and $2$ being chosen much more frequently than the midpoint between slits $2$ and $3$. The left side of the screen was also favored for peaks in correspondence with slits $1$ and $2$ and for the accumulation points at the ends of the screen, confirming a general bias favoring the left side of the screen. The origin of this asymmetry remains unclear. It may be attributed to the fact that participants are from a left-to-right reading culture, or it could be due to the population’s asymmetry between left-handed and right-handed individuals, but these are hypotheses that would require validation.
@@
When comparing our graph with those typically obtained in physics laboratories, the most striking difference is perhaps the fact that the central fringe is not the one of maximum intensity. A possible overall explanation for the existence of the two main peaks between slits $1$ and $2$, and between slits $2$ and $3$, for the secondary peak in correspondence with slit $3$, and for the two regions of accumulation of lesser intensity at the extremes of the screen, may be the following. When choosing a point $x$ on the screen, which should maximize the doubt about the slit from which the entity emerged, a participant will tend to associate with each slit a probability $\rho_{i}(x)$, $i=1,2,3$, inversely proportional to the distance $L_{i}$ between the slit and the point $x$ in question. In other words, the shorter the distance between the slit and point $x$, the more likely it is that the entity that produced that point emerged from that slit. More specifically, following Figure 11, we can write:
$L_{1}=\sqrt{L^{2}+\left(x+d\right)^{2}},\quad L_{2}=\sqrt{L^{2}+x^{2}},\quad L_{3}=\sqrt{L^{2}+\left(x-d\right)^{2}}$ (9)
$\rho_{i}(x)={A\over L_{i}},\quad A={L_{1}L_{2}L_{3}\over L_{1}L_{2}+L_{2}L_{3}+L_{1}L_{3}}$ (10)
Considering then the maximum $\rho(x)={\rm max}\{\rho_{1}(x),\rho_{2}(x),\rho_{3}(x)\}$, a participant trying to identify a point $x$ on the screen that maximizes the doubt about the slit from which the entity emerged will intuitively identify a position $x$ that minimizes $\rho(x)$. So, if $\overline{\rho}(x)=1-\rho(x)$ and $B=\sum_{x}\overline{\rho}(x)$, we can define $P_{\rm dis}(x)=B^{-1}\overline{\rho}(x)$ as the probability for selecting a point $x$ when using a reasoning based on competing distances. In Figure 7, we have plotted $P_{\rm dis}(x)$ together with $P(x|123)$. We see that $P_{\rm dis}(x)$ has its two main peaks precisely between slits $1$ and $2$, and between slits $2$ and $3$. This would explain the participants’ choice to favor those positions. Furthermore, the value of $P_{\rm dis}(x)$ is higher between the two slits than outside them, and this would explain the secondary choice of the central position. Finally, $P_{\rm dis}(x)$ shows accumulation points at the ends of the screen, which is also in line with the participants’ choices.
@@
Figure 7: (color online) Plot of the triple-slit probability $P(x|123)$ (red color) together with the probability $P_{\rm dis}(x)$ (black color) describing a reasoning based on the evaluation of distances between slits and impact points. To make the variations of $P_{\rm dis}(x)$ easier to read, it has been drawn using a distance $L=3$ (instead of $L=12$). As $L$ decreases (near-field limit), $P_{\rm dis}(x)$ increases at the edges of the screen, exceeding its value at the center.
It is interesting to ask whether it is possible to associate an effective wavelength that can account, at least approximately, for the observed patterns. For this, it is important to firstly note that the widths of our peaks depend on the scores attributed to the sequential choices of the participants. The data reported in Table 1 correspond to the choice of a linear scale: the first choice receive $7$ points, the second $6$ points, and so on, with the seventh receiving $1$ point, for a total of $28$ points per participant. However, we could have chosen to assign the same score to all seven sequential choices (uniform scale), in which case all seven choices receive $1$ point, for a total of $7$ points per participant. Another possibility is to use an exponential scale, with the first choice receiving $64$ points, the second $32$ points, and so on, with the seventh receiving $1$ point, for a total of $127$ points per participant. Finally, it is possible to consider the situation closest to what happens in a physics laboratory, using only the first choice of each participant, hence $1$ point per participant. In Figure 8, data relative to these four different choices of scales for the score assignment are compared, for the single-slit probabilities $P(x|1)$ and the three-slit probabilities $P(x|123)$.
Figure 8: A comparison of the data when changing the score assignment applied during analysis. In the left and right columns, the $P(x|1)$ and $P(x|123)$ probabilities are graphically represented, respectively. The first line corresponds to a uniform assignment (same score to all seven choices), the second line to a linearly decreasing assignment, the third line to an exponentially decreasing assignment, and finally the graphs of the last line only take into account participants’ best (first) choice. We observe that, as more weight is given to the first choices, the graphs become more sharply peaked.
To estimate the value of an effective wavelength $\lambda$ involved in the formation of the single-slit peaks, we have to use the graphs where only single answers are considered, otherwise the width of the peaks arbitrarily depends on the score assignment applied during analysis. We see in Figure 8 (fourth line) that peaks have a width of approximately $\Delta x\approx 3$. If we use the Fraunhofer diffraction formula (see Appendix A) $\Delta x\approx 2{\lambda L\over a}$, where $a$ is the width of the slit, since we have $a=1$ and $L=12$, the effective wavelength is $\lambda\approx a{\Delta x\over 2L}\approx{3\over 24}\approx 0.125$. On the other hand, If we apply the formula (see Appendix A) $\lambda\approx{xd\over L}$, giving the approximate distance $x$ at which a second maximum (the first being that of the central fringe) for the probability forms in the far field approximation, [^2] However, as we will discuss later in this article, this approximation does not adequately describe our experiment, given that the nullity of Sorkin’s parameter is strongly violated. considering (see Figure 6) $d=6$ and $x\approx{d\over 2}=3$, we find an effective wavelength $\lambda\approx{d^{2}\over 2L}=1.5$, which is clearly not compatible with the previously determined $\lambda\approx 0.125$ describing an effective diffraction mechanism. This means that there is no way to associate a single effective wavelength to our cognitive experiments, and this would be an important difference with the situation in physics laboratories (see also the discussion in Section 6).
Let us now consider the normalized Sorkin parameter $\kappa(x|123)$. Its maximum value is for $x=16$, i.e., in correspondence with slit-$1$: $\kappa(16|123)=0.549$; see Table 1, and the right graph of Figure 9. In the left graph of Figure 9, we have also plotted $\kappa(x|123)$ for the different score assignments (see also Figure 8). The maximum values obtained, for the normalized Sorkin parameter, are as follows: $0.508$ (uniform), $0.549$ (linear), $0.741$ (exponential) and $0.855$ (single impact).
Figure 9: (color online) Right: Plot of the normalized Sorkin parameter $\kappa(x|123)$ (blue color). The probability $P(x|123)$, multiplied by a factor of $10$, is also plotted (red color), to give a clearer understanding of the modulation in the plot. A continuous approximation of the discrete functions was used to enhance interpretability in the visualization. Left: Plots of the normalized Sorkin parameter $\kappa(x|123)$ for different choices of the scores attributed to the participants’ choices (uniform, linear, exponential and single impact); see Figure 8).
@@
We can observe that the values obtained are much higher than those typically obtained in physics laboratories. This reveals the presence of significant deviations of the probability $P(x|123)$ both from the classical situation and from the idealized quantum situation, when the principle of superposition can be applied naively. In other words, in their decision-making, participants were also guided by considerations that led to an overall assessment of the situation, taking into account the three slits together in their reasoning, rather than two at a time. This generated significant third-order interference-like contributions, which in physics can only be obtained in near-field situations, i.e., when the detection screen is not located far from the slits and/or the slits are very close together (distances on the order of the wavelength or less), so that non-radiative modes, such as evanescent waves, can easily mediate coupling between the slits and the detector, and/or between the slits.
It could be said that since the human cognitive instrument apparently uses different effective wavelengths at the same time, and wavelengths determine the scale of near-field dominance, in our double and triple-slit cognitive experiments we are always in regimes where near-field effects are important, hence the strong violation of the nullity of Sorkin parameter.